# Kermut: Composite kernel regression for protein variant effects

### v2

@@ Abstract
Reliable prediction of protein variant effects is crucial for both protein optimization and for advancing biological understanding. For practical use in protein engineering, it is important that we can also provide reliable uncertainty estimates for our predictions, and while prediction accuracy has seen much progress in recent years, uncertainty metrics are rarely reported. We here provide a Gaussian process regression model, Kermut, with a novel composite kernel for modelling mutation similarity, which obtains state-of-the-art performance for protein variant effect prediction while also offering estimates of uncertainty through its posterior. An analysis of the quality of the uncertainty estimates demonstrates that our model provides meaningful levels of overall calibration, but that instance-specific uncertainty calibration remains more challenging. We hope that this will encourage future work in this promising direction.
@@ 1 Introduction
Accurately predicting protein variant effects is crucial for both advancing biological understanding and for engineering and optimizing proteins towards specific traits. Recently, much progress has been made in the field as a result of advances in machine learning-driven modeling [1, 2, 3], data availability [4], and relevant benchmarks [5, 6].
While prediction accuracy has received considerable attention, the ability to quantify the uncertainties of predictions has been less intensely explored. This is of immediate practical consequence. One of the main purposes of protein variant effect prediction is as an aide for protein engineering and design, to propose promising candidates for subsequent experimental characterization. For this purpose, it is essential that we can quantify, on an instance-to-instance basis, how trustworthy our predictions are. Specifically, in a Bayesian optimization setting, most choices of acquisition function actively rely on predicted uncertainties to guide the optimization, where well-calibrated uncertainties have been show to correlate with optimization performance [7].
Our goal with this paper is to start a discussion on the quality of the estimated uncertainties of protein property predictions. Gaussian Process (GP) regression is a standard choice for uncertainty quantification due to the closed form expression for the posterior. We therefore first ask the question whether state-of-the-art performance can be obtained within the GP framework. We propose a composite kernel that comfortably achieves this goal, and subsequently investigate the quality of the uncertainty estimates from such a model. Our results show that while standard approaches like reliability diagrams give the impression of good levels of calibration, the quantification of per-instance uncertainties is more challenging. We make our model available as a baseline and encourage the community to place greater emphasis on uncertainty quantification in this important domain. Our contributions can be summarized as follow:
  • We introduce Kermut, a Gaussian process with a novel composite kernel for modelling mutation similarity, leveraging signals from pretrained sequence and structure models;
  • We evaluate this model on the comprehensive ProteinGym substitution benchmark and show that it is able to reach state-of-the-art performance in protein variant effect prediction, outperforming recently proposed deep learning methods in this domain;
  • We provide a thorough calibration analysis and show that Kermut provides well-calibrated uncertainties in most cases;
@@ 2.1 Protein property prediction
Recently, unsupervised approaches have gained significant momentum where models trained in a self-supervised fashion have shown impressive results for zero-shot estimates of protein fitness and variant effects, where the effect of introducing mutations to a reference protein are gauged [3, 8, 9, 10].
@@
Supervised learning remains a crucial method of utilizing experimental data to predict protein fitness. This is particularly valuable when the trait of interest correlates poorly with the evolutionary signals that unsupervised models capture during training or if multiple traits are considered. Supervised protein fitness prediction using machine learning has been explored in detail in [11], where a comprehensive overview can be found. A major change since then is the use of transfer learning via embeddings extracted from self-supervised models, such as protein language models like ESM-2 [12] and ProtTrans [13]. In [14], the authors propose to augment a one-hot encoding of the aligned amino acid sequence by concatenating it with a zero-shot score for improved predictions. This was further expanded upon with ProteinNPT [15], where sequences embedded with the MSA Transformer [16] and zero-shot scores were fed to a transformer architecture for state-of-the-art property prediction.
@@
Considerable progress has been made in defining meaningful and comprehensive benchmarks to reliably measure and compare model performance in both unsupervised and supervised protein fitness prediction settings. The FLIP benchmark [5] introduced three supervised predictions tasks ranging from local to global fitness prediction, where each task in turn was divided into clearly defined splits. The supervised benchmarks often view fitness prediction through a particular lens. Where FLIP targeted problems of interest to protein engineering; TAPE [17] evaluated transfer learning abilities; PEER [18] focused on sequence understanding; ATOM3D [19] considered a structure-based approach; FLOP [20] targeted wild type proteins; and ProteinGym focused exclusively on variant-effect prediction [10]. The ProteinGym benchmark was recently expanded to encompass more than 200 standardized datasets in both zero-shot and supervised settings, including substitutions, insertions, deletions, and curated clinical datasets [10, 6].
2.2 Kernel methods
@@
Kernel methods have seen much use for protein property prediction. Sequence-based string kernels operating directly on the protein amino acid sequences are one such example, where, e.g., matching k-mers at different ks quantify covariance. This has been used with support vector machines to predict protein homology [21, 22]. In [23], a kernel leveraging the tertiary structure for a protein family represented as a residue-residue contact map, was used to predict various protein properties such as enzymatic activity and binding affinity. In [24], a graph-based weighted decomposition kernel operating on tertiary protein structure in conjunction with substitution matrices was used to predict experimental $\Delta\Delta$G values of protein variants.
@@ 2.3 Local structural environments
Much work has been done to solve the inverse-folding problem, where the most probable amino acid sequence to fold into a given protein backbone structure is predicted [25, 26, 27, 28]. Inverse-folding models are trained on large datasets of proteins structures and model the physicochemical and evolutionary constraints of sites in a protein conditioned on their structural context. These will form the basis of the the structural featurization in our work. Local structural environments were previously used to model mutational preferences for protein engineering [29], however not in a Gaussian process framework as we propose here. Structural environments have also been used for surface-level fingerprinting which was subsequently used to model protein-protein interaction sites and for de novo design of protein interactions [30, 31].
@@ 3.1 Preliminaries
Our goal is to predict the outcome of an assay measured on a protein, represented by its amino acid sequence $\mathbf{x}$ of length $L$. We will assume that we have a dataset of N such sequences available, and that these are of equal length and structure, such that we can meaningfully refer to the effect at specific positions (sites) in the protein. In protein engineering, we typically consider modifications relative to an initial wild type sequence, $\mathbf{x}_{\text{WT}}$. We will assume that the 3D structure for the initial sequence, $s$, is available (either experimentally determined or provided by a structure predictor like AlphaFold [32]). Lastly, for variant $\mathbf{x}$ with mutations at sites $M\subseteq\{1,...,L\}$, let $\mathbf{x}^{m}$ denote the variant which has the same mutation as $\mathbf{x}$ at site $m$ for $m\in M$ and otherwise is equal to $\mathbf{x}_{\text{WT}}$.
@@ 3.2 Gaussian processes
To predict protein variant effects we rely on Gaussian process regression, which we shall now briefly introduce. For a comprehensive overview, see [33], which this section is based on.
@@
Gaussian processes are stochastic processes providing a powerful framework for modelling distributions over functions. The Gaussian process framework allows us to not only make predictions but also to quantify the uncertainty associated with each prediction. A Gaussian process is entirely specified by its mean and covariance functions, $m(\mathbf{x})$ and $k(\mathbf{x},\mathbf{x^{\prime}})$. We assume that the covariance matrix, $K$, of the outputs $\{y_{1},...,y_{N}\}$ can be parameterized by a function of their inputs $\{\mathbf{x}_{1},...,\mathbf{x}_{N}\}$. The parameterization is defined by the kernel, $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ yielding $K$ such that $K_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$. For $k$ to be valid kernel, $K$ needs to be symmetric and positive semidefinite.
@@ 3.3 Kermut
In [34], the authors showed that a simple linear model operating on one-hot encoded mutations is sufficient to accurately predict mutation effects given enough data. Starting with a linear model, we explore how to improve predictive performance while staying within the realm of models with inherent ways of providing uncertainty estimates.
We begin by assuming that the variant effect relative to wild type protein $\mathbf{x}_{\text{WT}}$ is additive in terms of amino acids for each site. We can then write $f(\mathbf{x})=\sum_{m\in M}f(\mathbf{x}^{m}),$ where $f:\mathbb{R}^{19L}\rightarrow\mathbb{R}$. We furthermore assume that $f$ is a linear function: $f(\mathbf{x})=\mathbf{x}\theta$. For $N$ observations, we write $f(\mathbf{X})=\mathbf{X}\theta$, where $\mathbf{X}$ is a $N\times 19L$ design matrix and $\theta$ is a $19L$ dimensional effect vector, where the first 19 elements corresponds to the 19 possible mutations at site one, etc. To reduce the amount of data needed in the high-dimensional space of possible variants, we make a Bayesian interpretation of the function and restrict the flexibility of the model by specifying a prior belief over the weights that reflects prior biological knowledge. For example, we can use the prior $\theta\sim\mathcal{N}(0,\Sigma_{\theta})$ to specify that mutations on the same site are expected to have similar effects, with $\Sigma_{\theta ij}=1$ if $i=j$, $\Sigma_{\theta ij}=c\in[0,1]$ if $i$ and $j$ corresponds to mutations on the same site, and $\Sigma_{\theta ij}=0$ otherwise. Additional prior knowledge can be imposed by adjusting $\Sigma_{\theta}$ accordingly. We can also use a feature map $\phi(\mathbf{x})$ as input to the linear function: $f(\mathbf{X})=\mathbf{\Phi(X)}\theta$, where $\mathbf{\Phi(X)}=[\phi(\mathbf{x}_{1}),...,\phi(\mathbf{x}_{n})]^{T}$. Having chosen a feature map $\phi$ and a prior covariance matrix $\Sigma_{\theta}$, then $f(\mathbf{X})\sim\mathcal{N}(0,\mathbf{\Phi(X)}\Sigma_{\theta}\mathbf{\Phi(X)}^{T})$. We may reformulate this distribution over function values as a Gaussian process with covariance function $K(\mathbf{X})=\mathbf{\Phi(X)}\Sigma_{\theta}\mathbf{\Phi(X)}^{T}$ and zero mean. By using the GP framework, a large set of flexible covariance functions to reflect prior knowledge are available. For the problem at hand, the prior specifies the covariance between function values of the protein variants.
@@
Arguably, a straightforward strategy for constructing a kernel for protein regression is to use a squared exponential kernel directly on a sequence embedding extracted from a pretrained model [34]. We use the ESM-2 protein language model [12], and perform mean-pooling across the length dimension, yielding $\mathbf{z}=f_{1}(\mathbf{x})$, where $f_{1}$ produces mean-pooled embeddings, $\mathbf{z}\in\mathbb{R}^{1280}$, of sequence $\mathbf{x}$. We thus model the covariance between these representations as
@@
It has been shown that local structural dependencies are useful for determining mutation preferences [29]. As $k_{\text{seq}}$ only relies on embedded sequences, we hypothesize that incorporating information about the local structural environments will lead to improved covariance estimates. To this end, we define a structure kernel, $k_{\text{struct}}$, which models mutation similarity given the local environments of mutated sites. A schematic of how the structure kernel models covariances can be seen in Figure 1.
Specifically, we hypothesize that for a given site in the protein, the distribution over amino acids given by a structure-conditioned inverse folding model will reflect the effect of a mutation at that site. We consider such an amino acid distribution a representation of the local environment for that site as it reflects the various physicochemical and evolutionary constraints that the site is subject to. We thus presume that two sites with similar local environments will behave similarly if mutated. For instance, mutations at buried sites in the hydrophobic core of the protein will generally correlate more with each other than with surface-level mutations.
@@
GP Kermut 0.611 0.634 0.747 0.664 0.699 0.651 0.409 0.586
For single mutant variants we quantify site similarity using the Hellinger kernel $k_{H}(\mathbf{x},\mathbf{x^{\prime}})=\exp\left(-\gamma_{1}d_{H}(f_{\text{IF}}(\mathbf{x}),f_{\text{IF}}(\mathbf{x}^{\prime}))\right)$, with $\gamma_{1}>0$ [35], where $d_{H}$ is the Hellinger distance (see Appendix A.1). The function $f_{\text{IF}}:\mathcal{X}^{1}\rightarrow[0,1]^{20}$ takes a single-mutant sequence, $\mathbf{x}$, as input and returns a probability distribution over the 20 naturally occurring amino acids at the mutated site in $\mathbf{x}$ given by an inverse folding model. The Hellinger kernel will assign maximum covariance when two sites are identical, preventing the comparison of different mutations at the same site.
To increase flexibility and to allow intra-site comparisons, we introduce a kernel operating on the specific mutation likelihoods. We hypothesize that two variants with mutations on sites that are close in terms of the Hellinger distance will correlate further if the log-probabilities of the specific amino acids on the mutated sites are similar (i.e., the probability of the amino acid that we mutate to is similar at the two sites). We incorporate this by defining $k_{p}(\mathbf{x},\mathbf{x^{\prime}})=k_{\text{exp}}(f_{\text{IF}_{1}}(\mathbf{x}),f_{\text{IF}_{1}}(\mathbf{x}^{\prime}))$, where $f_{\text{IF}_{1}}:\mathcal{X}^{1}\rightarrow[0,1]$ takes a single-mutant sequence, $\mathbf{x}$, as input and returns the log-probability (given by an inverse folding model) of the observed mutation, and where $k_{\text{exp}}$ is the exponential kernel.
Finally, we hypothesize that the effect of two mutations correlate further if the sites are close in physical space. Hence, we multiply the kernel with an exponential kernel on the Euclidean distance between sites: $k_{d}(\mathbf{x},\mathbf{x^{\prime}})=\exp\left(-\gamma_{2}d_{e}(s_{i},s_{j})\right)$. Thereby, the closer two sites are, the more similar – and thus comparable – their local environments will be.
For single-mutant variants we now have the following kernel:
where the kernel has been scaled by a non-negative scalar, $\lambda>0$. To generalize the kernel to multiple mutations, we simply sum over all pairs of sites differing at $\mathbf{x}$ and $\mathbf{x}^{\prime}$:
We add the structure and sequence kernels resulting in a final formulation to model covariances between mutations:
Additional details on both sequence and structure kernels, including a proof of the validity of the structure kernel as well as implementation details can be found in Appendices A and B.

### v3

@@ Abstract
Reliable prediction of protein variant effects is crucial for both protein optimization and for advancing biological understanding. For practical use in protein engineering, it is important that we can also provide reliable uncertainty estimates for our predictions, and while prediction accuracy has seen much progress in recent years, uncertainty metrics are rarely reported. We here provide a Gaussian process regression model, Kermut, with a novel composite kernel for modeling mutation similarity, which obtains state-of-the-art performance for supervised protein variant effect prediction while also offering estimates of uncertainty through its posterior. An analysis of the quality of the uncertainty estimates demonstrates that our model provides meaningful levels of overall calibration, but that instance-specific uncertainty calibration remains more challenging.
@@ 1 Introduction
Accurately predicting protein variant effects is crucial for both advancing biological understanding and for engineering and optimizing proteins towards specific traits. Recently, much progress has been made in the field as a result of advances in machine learning-driven modeling [1, 2, 3], data availability [4, 5], and relevant benchmarks [6, 7].
While prediction accuracy has received considerable attention, the ability to quantify the uncertainties of predictions has been less intensely explored. This is of immediate practical consequence. One of the main purposes of protein variant effect prediction is as an aid for protein engineering and design, to propose promising candidates for subsequent experimental characterization. For this purpose, it is essential that we can quantify, on an instance-to-instance basis, how trustworthy our predictions are. Specifically, in a Bayesian optimization setting, most choices of acquisition function actively rely on predicted uncertainties to guide the optimization, and well-calibrated uncertainties have been shown to correlate with optimization performance [8].
Our goal with this paper is to start a discussion on the quality of the estimated uncertainties of supervised protein variant effect prediction. Gaussian Processes (GP) are a standard choice for uncertainty quantification due to the closed form expression of the posterior. We therefore first ask the question whether state-of-the-art performance can be obtained within the GP framework. We propose a composite kernel that comfortably achieves this goal, and subsequently investigate the quality of the uncertainty estimates from such a model. Our results show that while standard approaches like reliability diagrams give the impression of good levels of calibration, the quantification of per-instance uncertainties is more challenging. We make our model available as a baseline and encourage the community to place greater emphasis on uncertainty quantification in this important domain. Our contributions can be summarized as follow:
  • We introduce Kermut, a Gaussian process with a novel composite kernel for modeling mutation similarity, leveraging signals from pretrained sequence and structure models;
  • We evaluate this model on the comprehensive ProteinGym substitution benchmark and show that it is able to reach state-of-the-art performance in supervised protein variant effect prediction, outperforming recently proposed deep learning methods in this domain;
  • We provide a thorough calibration analysis and show that while Kermut provides well-calibrated uncertainties overall, the calibratedness of instance-specific uncertainties remains challenging;
@@ 2.1 Protein property prediction
Recently, unsupervised approaches have gained significant momentum where models trained in a self-supervised fashion have shown impressive results for zero-shot estimates of protein fitness and variant effects relative to a reference protein [3, 9, 10, 11].
@@
Supervised learning is a crucial method of utilizing experimental data to predict protein fitness. This is particularly valuable when the trait of interest correlates poorly with the evolutionary signals that unsupervised models capture during training or if multiple traits are considered. Supervised protein fitness prediction using machine learning has been explored in detail in [12], where a comprehensive overview can be found. A common strategy is to employ transfer learning via embeddings extracted from self-supervised models [13, 14], an approach which increasingly relies on large pretrained language models such as ProtTrans [15], ESM-2 [16], and SaProt [17]. In [18], the authors propose to augment a one-hot encoding of the aligned amino acid sequence by concatenating it with a zero-shot score for improved predictions. This was further expanded upon with ProteinNPT [19], where sequences embedded with the MSA Transformer [20] and zero-shot scores were fed to a transformer architecture for state-of-the-art supervised variant effect prediction with generative capabilities.
@@
Considerable progress has been made in defining meaningful and comprehensive benchmarks to reliably measure and compare model performance in both unsupervised and supervised protein fitness prediction settings. The FLIP benchmark [6] introduced three supervised predictions tasks ranging from local to global fitness prediction, where each task in turn was divided into clearly defined splits. The supervised benchmarks often view fitness prediction through a particular lens. Where FLIP targeted problems of interest to protein engineering; TAPE [21] evaluated transfer learning abilities; PEER [22] focused on sequence understanding; ATOM3D [23] considered a structure-based approach; FLOP [24] targeted wild type proteins; and ProteinGym focused exclusively on variant effect prediction [11]. The ProteinGym benchmark was recently expanded to encompass more than 200 standardized datasets in both zero-shot and supervised settings, including substitutions, insertions, deletions, and curated clinical datasets [11, 7].
2.2 Kernel methods for protein sequences
@@
Kernel methods have seen much use for protein modeling and protein property prediction. Sequence-based string kernels operating directly on the protein amino acid sequences are one such example, where, e.g., matching k-mers at different ks quantify covariance. This has been used with support vector machines to predict protein homology [25, 26]. Another example is sub-sequence string kernels, which in [27] is used in a Gaussian process for a Bayesian optimization procedure. In [28], string kernels were combined with predicted physicochemical properties to improve accuracy in the prediction of MHC-binding peptides and in protein fold classification. In [29], a kernel leveraging the tertiary structure for a protein family represented as a residue-residue contact map was used to predict various protein properties such as enzymatic activity and binding affinity. In [30], Gaussian process regression (GPR) was used to successfully identify promising enzyme sequences which were subsequently synthesized showing increased activity. In [31], the authors provide a comprehensive study of kernels on biological sequences which includes a thorough review of the literature as well as both theoretical, simulated, and in-silico results.
Most similar to our work is mGPfusion [32], in which a weighted decomposition kernel was defined which operated on the local tertiary protein structure in conjunction with a number of substitution matrices. Simulated stability data for all possible single mutations were obtained via Rosetta [33], which was then fused with experimental data for accurate $\Delta\Delta$G predictions of single- and multi-mutant variants via GPR, thus incorporating both sequence, structure, and a biophysical simulator. In contrast to our approach, the mGPfusion-model does not leverage pretrained models, but instead relies on substitution matrices for its evolutionary signal. A more recent example of kernel-based methods yielding highly competitive results is xGPR [34], in which Gaussian processes with custom kernels show high performance when trained on protein language model embeddings, similarly to the sequence kernel in our work (see Section 3.3). Where xGPR introduces a set of novel random feature-approximated kernels with linear-scaling, Kermut instead uses the squared exponential kernel for sequence modeling while additionally modeling local structural environments. The models in xGPR were shown to provide both high accuracy and well-calibrated uncertainty estimation on the FLIP and TAPE benchmarks.
2.3 Uncertainty quantification and calibration
Uncertainty quantification (UQ) for protein property prediction continues to be a promising area of research with immediate practical consequences. In [35], residual networks were used to model both epistemic and aleatoric uncertainty for peptide selection. In [36], GPR on MLP-residuals from biLSTM embeddings was used to successfully guide in-silico experimental design of kinase binders and fluorescent proteins. The authors of [37] augmented a Bayesian neural network by placing biophysical priors over the mean function by directly using Rosetta energy scores, whereby the model would revert to the biophysical prior when the epistemic uncertainty was large. This was used to predict fluorescence, binding, and solubility for drug-like molecules. In [38], state-of-the-art performance on protein-protein interactions was achieved by using a spectral-normalized neural Gaussian process [39] with an uncertainty-aware transformer-based architecture working on ESM-2 embeddings.
In [40], a framework for evaluating the epistemic uncertainty of deep learning models using confidence interval-based metrics was introduced, while [41] conducted a thorough analysis of uncertainty quantification methods for molecular property prediction. Here, the importance of supplementing confidence-based calibration with error-based calibration as introduced in [42] was highlighted, whereby the predicted uncertainties are connected directly to the expected error for a more nuanced calibration analysis. We evaluate our model using confidence-based calibration as well as error-based calibration following the guidelines in [41]. In [43], the authors conducted a systematic comparison of UQ methods on molecular property regression tasks, while [44] investigated calibratedness of regression models for material property prediction. In [45], the above approaches were expanded to protein property prediction tasks where the FLIP [6] benchmark was examined, while [46] benchmarked a number of UQ methods for molecular representation models. In [47], the authors developed an active learning approach for partial charge prediction of metal-organic frameworks via Monte Carlo dropout [48] while achieving decent calibration. In [49], a systematic analysis of protein regression models was conducted where well-calibrated uncertainties were observed for a range of input representations.
@@ 2.4 Local structural environments
Much work has been done to solve the inverse-folding problem, where the most probable amino acid sequence to fold into a given protein backbone structure is predicted [50, 51, 52, 53, 54, 55, 56]. Inverse-folding models are trained on large datasets of protein structures and model the physicochemical and evolutionary constraints of sites in a protein conditioned on their structural contexts. These will form the basis of the structural featurization in our work. Local structural environments have previously been used for protein modeling. In [57], a 3D CNN was used to predict amino acid preferences giving rise to novel substitution matrices. In [58] and [59], surface-level fingerprinting given structural environments was used to model protein-protein interaction sites and for de novo design of protein interactions. In [60], chemical microenvironments were used to identify potentially beneficial mutations, while [61] used a similar approach, where they investigated the volume of the local environments and observed that the first contact shell delivered the primary signal thus emphasizing the importance of locality. In [62], a composition Hellinger distance metric based on the chemical composition of local residue environments was developed and used for a range of structure-related experiments. Recently, local structural environments were used to model mutational preferences for protein engineering tasks [63], however not in a Gaussian process framework as we propose here.
@@ 3.1 Preliminaries
We want to predict the outcome of an assay measured on a protein, represented by its amino acid sequence $\mathbf{x}$ of length $L$. We will assume that we have a dataset of N such sequences available, and that these are of equal length and structure, such that we can meaningfully refer to the effect at specific positions (sites) in the protein. In protein engineering, we typically consider modifications relative to an initial wild type sequence, $\mathbf{x}_{\text{WT}}$. We will assume that the 3D structure for the initial sequence, $s$, is available (either experimentally determined or provided by a structure predictor like AlphaFold [64]). Lastly, for variant $\mathbf{x}$ with mutations at sites $M\subseteq\{1,...,L\}$, let $\mathbf{x}^{m}$ denote the variant which has the same mutation as $\mathbf{x}$ at site $m$ for $m\in M$ and otherwise is equal to $\mathbf{x}_{\text{WT}}$.
@@ 3.2 Gaussian processes
To predict protein variant effects, we rely on Gaussian process regression, which we shall now briefly introduce. For a comprehensive overview, see [65], which this section is based on.
@@
Gaussian processes are stochastic processes providing a powerful framework for modeling distributions over functions. The Gaussian process framework allows us to not only make predictions but also to quantify the uncertainty associated with each prediction. A Gaussian process is entirely specified by its mean and covariance functions, $m(\mathbf{x})$ and $k(\mathbf{x},\mathbf{x^{\prime}})$. We assume that the covariance matrix, $K$, of the outputs $\{y_{1},...,y_{N}\}$ can be parameterized by a function of their inputs $\{\mathbf{x}_{1},...,\mathbf{x}_{N}\}$. The parameterization is defined by the kernel, $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ yielding $K$ such that $K_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$. For $k$ to be a valid kernel, $K$ needs to be symmetric and positive semidefinite.
@@ 3.3 Kermut
It has been shown that local structural dependencies are useful for determining mutation preferences [57, 60, 61, 63]. We therefore hypothesize that constructing a composite kernel with components incorporating information about the local structural environments of residues will be able to model protein variant effects. To this end, we define a structure kernel, $k_{\text{struct}}$, which models mutation similarity given the local environments of mutated sites. A schematic of how the structure kernel models covariances can be seen in Figure 1. In the following we shall define $k_{\text{struct}}^{1}$, a structure kernel that operating on single-mutant variants. Subsequently, we shall extend it to multi-mutant variants resulting in the structure kernel, $k_{\text{struct}}$.
We hypothesize that for a given site in a protein, the distribution over amino acids given by a structure-conditioned inverse folding model will reflect the effect of a mutation at that site. We consider such an amino acid distribution a representation of the local environment for that site as it reflects the various physicochemical and evolutionary constraints that the site is subject to. We thus presume that two sites with similar local environments will behave similarly if mutated. For instance, mutations at buried sites in the hydrophobic core of the protein will generally correlate more with each other than with surface-level mutations.
For single mutant variants we quantify site similarity using the Hellinger kernel $k_{H}(\mathbf{x},\mathbf{x^{\prime}})=\exp\left(-\gamma_{1}d_{H}(f_{\text{IF}}(\mathbf{x}),f_{\text{IF}}(\mathbf{x}^{\prime}))\right)$, with $\gamma_{1}>0$ [66], where $d_{H}$ is the Hellinger distance (see Appendix B.1). The function $f_{\text{IF}}:\mathcal{X}^{1}\rightarrow[0,1]^{20}$ takes a single-mutant sequence, $\mathbf{x}$, as input and returns a probability distribution over the 20 naturally occurring amino acids at the mutated site in $\mathbf{x}$ given by the inverse folding model. The Hellinger kernel will assign maximum covariance when two sites are identical. This however means that $k_{H}$ is incapable of distinguishing between different mutations at the same site since $d_{H}(\mathbf{x},\mathbf{x}^{\prime})=1$, when $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are mutated at the same site.
To increase flexibility and to allow intra-site comparisons, we introduce a kernel operating on the specific mutation likelihoods. We hypothesize that two variants with mutations on sites that are close in terms of the Hellinger distance will correlate further if the log-probabilities of the specific amino acids on the mutated sites are similar (i.e., the probability of the amino acid that we mutate to is similar at the two sites). We incorporate this by defining $k_{p}(\mathbf{x},\mathbf{x^{\prime}})=k_{\text{exp}}(f_{\text{IF}_{1}}(\mathbf{x}),f_{\text{IF}_{1}}(\mathbf{x}^{\prime}))=\text{exp}(-\gamma_{2}||f_{\text{IF}_{1}}(\mathbf{x})-f_{\text{IF}_{1}}(\mathbf{x}^{\prime})||)$, where $f_{\text{IF}_{1}}:\mathcal{X}^{1}\rightarrow[0,1]$ takes a single-mutant sequence, $\mathbf{x}$, as input and returns the log-probability (given by an inverse folding model) of the observed mutation, and where $k_{\text{exp}}$ is the exponential kernel.
Finally, we hypothesize that the effect of two mutations correlate further if the sites are close in physical space. Hence, we multiply the kernel with an exponential kernel on the Euclidean distance between sites: $k_{d}(\mathbf{x},\mathbf{x^{\prime}})=\exp\left(-\gamma_{3}d_{e}(s_{i},s_{j})\right)$. Thereby, the closer two sites are physically, the more similar – and thus comparable – their local environments will be.
Taking the product of these kernel components, we get the following kernel for single-mutant variants, which assigns high covariance when two single mutant variants have mutations that have similar environments, are physically close, and have similar mutation likelihoods:
where the kernel has been scaled by a non-negative scalar, $\lambda>0$.
In [63], the authors showed that a simple linear model operating on one-hot encoded mutations is sufficient to accurately predict mutation effects given sufficient data. Thus, we generalize the kernel to multiple mutations by summing over all pairs of sites differing at $\mathbf{x}$ and $\mathbf{x}^{\prime}$:
@@
The structure kernel models mutations linearly and cannot capture epistatic effects. We propose to add epistatic signals through a sequence kernel. Drawing on the rich literature for modeling protein sequences with Gaussian processes on embeddings [34, 45, 49, 67], we use a squared exponential kernel which operates on sequence embeddings from a pretrained model. We use the 650M parameter ESM-2 protein language model [16], and perform mean-pooling across the length dimension, yielding $\mathbf{z}=f_{1}(\mathbf{x})$, where $f_{1}$ produces mean-pooled embeddings, $\mathbf{z}\in\mathbb{R}^{1280}$, of sequence $\mathbf{x}$. We thus model the covariance between these representations as
@@
We choose to add and weigh the structure and sequence kernels resulting in our final kernel formulation, whereby the model can leverage either structure or sequence similarities, depending on the presence and strength of each signal as determined through hyperparameter optimization:
Additional details on both sequence and structure kernels, including a proof of the validity of the structure kernel, implementation details, computational complexity details, and an automatic model selection procedure can be found in Appendices B and C.
@@
GP Kermut 0.610 0.633 0.744 0.662 0.699 0.652 0.414 0.589